### v1

# Yunjue Agent Tech Report: A Fully Reproducible, Zero-Start In-Situ Self-Evolving Agent System for Open-Ended Tasks

Abstract
Conventional agent systems often struggle in open-ended environments where task distributions continuously drift and external supervision is scarce. Their reliance on static toolsets or offline training lags behind these dynamics, leaving the system’s capability boundaries rigid and unknown. To address this, we propose the In-Situ Self-Evolving paradigm. This approach treats sequential task interactions as a continuous stream of experience, enabling the system to distill short-term execution feedback into long-term, reusable capabilities without access to ground-truth labels. Within this framework, we identify tool evolution as the critical pathway for capability expansion, which provides verifiable, binary feedback signals. Within this framework, we develop Yunjue Agent, a system that iteratively synthesizes, optimizes, and reuses tools to navigate emerging challenges. To optimize evolutionary efficiency, we further introduce a Parallel Batch Evolution strategy. Empirical evaluations across five diverse benchmarks under a zero-start setting demonstrate significant performance gains over proprietary baselines. Additionally, complementary warm-start evaluations confirm that the accumulated general knowledge can be seamlessly transferred to novel domains. Finally, we propose a novel metric to monitor evolution convergence, serving as a function analogous to training loss in conventional optimization. We open-source our codebase, system traces, and evolved tools to facilitate future research in resilient, self-evolving intelligence.
Machine Learning, ICML
1 Introduction
The rapid ascendancy of Large Language Models (LLMs) has catalyzed a paradigm shift towards Artificial General Intelligence (AGI) (Guo et al., 2024). While this wave has birthed versatile agents, a critical dichotomy remains: the most capable systems are increasingly dominated by closed-source paradigms (Phan et al., 2025; Gupta et al., 2025). These proprietary models often rely on opaque tools and black-box APIs, creating a barrier for verifiable scientific inquiry. Conversely, open-source alternatives, while accessible, frequently lag in performance (Manchanda et al., 2024; Choi and Chang, 2025), as their reliance on static, manually crafted heuristics renders them unable to rival the robustness of industrial proprietary systems (Jin et al., 2024). We posit that for open-source intelligence to bridge this gap, it must move beyond static imitation of closed models and leverage a dynamic advantage: In-Situ Self-Evolution, where the system becomes stronger through continuous usage.
We envision that true AGI demands an In-Situ Self-Evolving architecture. In this ideal state, an agent acts as a fully adaptive organism, continuously refining its entire cognitive stack—workflow, context and tool. Here, the workflow adapts planning strategies based on user preferences or task types (Ye et al., 2025); the context dynamically distills episodic experiences into wisdom (Tang et al., 2025); and the tools evolve autonomously to conquer novel execution boundaries (Qin et al., 2024).
Recently, the field has witnessed a surge in self-evolving agents (Tao et al., 2024; Gao et al., 2025) designed to enhance these components autonomously (Xi et al., 2025; Cai et al., 2025; Acikgoz et al., 2025). Progress has been made in workflow adaptation (Zhang et al., 2025a) and context management (Tang et al., 2025). However, a critical limitation persists: existing paradigms predominantly rely on offline training with explicit reward signals or remain confined to narrow, domain-specific boundaries.
In supervision-scarce, open-ended environments—where ground truth is absent and user feedback is sparse—this reliance on external supervision becomes a bottleneck. Evolving a workflow or aligning with user preferences often involves subjective or delayed rewards (e.g., “Did the user like this summary?”). In contrast, tool evolution offers a distinct advantage: the feedback signal is intrinsic and objective (Wang et al., 2024a). Regardless of the final task outcome, the functionality of a tool is verifiable via rigorous execution feedback—code either runs successfully or throws an exception. This robust signal allows for autonomous optimization even in the absence of human intervention, making tools the most logical starting point for in-situ evolution. Beyond the signal quality, the toolset also constitutes the fundamental operational prerequisite for any agent (Zaremba et al., 2021). Building on this insight, we advocate for a staged, foundational approach that prioritizes the evolution of tools.
To this end, we introduce Yunjue Agent, a framework that prioritizes the autonomous synthesis, validation, and convergence of tools under a zero-start paradigm. Mechanically, the system employs a multi-agent architecture (Qian et al., 2024), comprising a Manager, Tool Developer, and Executor to address queries dynamically. Instead of relying on a static library, the agent synthesizes bespoke Python primitives on-the-fly when existing capabilities are insufficient. Crucially, to prevent tool explosion and ensure robustness, we introduce a Parallel Batch Evolution strategy. This mechanism processes queries in batches, clustering and merging functionally similar tools via an “absorbing” process. This allows the system to distill transient, query-specific tool into a compact, convergent repository of generalized knowledge.
Our contributions are summarized as follows:
  1. 1.
    SOTA Performance via zero-start In-Situ Self-Evolving: We propose Yunjue Agent, a system that synthesizes, validates, and refines tools from scratch without prior training, with SOTA results on rigorous benchmarks.
  2. 2.
    Cross-domain transferability: We demonstrate that evolved tools exhibit robust transferability across disjoint domains, proving that the system distills transient interactions into generalized, convergent capabilities.
  3. 3.
    Evolution convergence metric: We introduce a quantitative metric to monitor evolutionary stability during inference, serving as a function analogous to “training loss” in conventional optimization.
  4. 4.
    Reproducibility and transparency: We open-source our codebase, complete interaction traces, and the library of evolved tools to provide a transparent foundation for future research.
2 In-situ self-evolving agents
An agent system can be formally defined as a tuple $\mathcal{M}=\langle\mathcal{W},\mathcal{C},\mathcal{T}\rangle$ of workflow, context and tools (Gao et al., 2025). In this system, the workflow $\mathcal{W}$ is typically structured as a directed graph $\mathcal{G}=(V,E)$, where nodes $V$ represent LLM-based agents and edges $E$ denote the flow of information. $\mathcal{C}$ is the set of contexts, where each context element can be a prompt template or a memory buffer encompassing dynamic conversation history. Finally, $\mathcal{T}$ denotes the set of tools available for the agents to execute specific tasks.
To enable agents to adapt to new environments like humans, recent work focuses on achieving this through self-evolving methods (Tao et al., 2024; Gao et al., 2025). However, these methods typically require an iterative training process (He et al., 2025; Chen et al., 2025a) to update certain components or are confined in specific areas (Xia et al., 2025; Jin et al., 2025), such as generating query-specific workflows (Zhang et al., 2025a; Ye et al., 2025), refining context $\mathcal{C}$ (e.g., prompt and memory) (Zhang et al., 2025a; Tang et al., 2025) or adjusting toolset (Wang et al., 2024b, 2023). In contrast, we aim to explore whether an agent system can adapt to the environment by continuously updating its components in-situ, where there is no access to external supervision signals. This motivation leads to a new paradigm of agentic evolution.
In-situ self-evolving of agents. Given a sequence of queries $\{x_{1},x_{2},\dots,x_{T}\}$, the agent evolves dynamically from $\mathcal{M}_{0}$ to $\mathcal{M}_{T}$ as it processes each. Specifically, after completing the $t$-th query $x_{t}$, the system updates its configuration for the next query, which can be formalized as a transition from $\mathcal{M}_{t-1}=\langle\mathcal{W}_{t-1},\mathcal{C}_{t-1},\mathcal{T}_{t-1}\rangle$ to $\mathcal{M}_{t}=\langle\mathcal{W}_{t},\mathcal{C}_{t},\mathcal{T}_{t}\rangle$, where any component of the tuple—the workflow structure, context, or toolset—may be modified based on internal feedback or experience gained from the previous interaction. Unlike self-evolving agents that maximize a specific objective function (e.g., accuracy) through a series of training steps (Gao et al., 2025), in-situ self evolving operates during the inference phase where ground truth information is unavailable.
To ensure a focused scope, we fix the workflow $\mathcal{W}$. Furthermore, given that Agent KB (Tang et al., 2025) has attempted to optimize memory through cross-domain experience to enhance agent performance, we also treat the context $\mathcal{C}$ as a fixed component [^1] Although the execution flow can be dynamic during inference due to branch control and model inputs are populated at runtime, the underlying workflow structure and prompt templates are pre-defined. Thus, we consider these configurations fixed., reducing the evolving system state to $\mathcal{M}_{t}=\langle\mathcal{W}_{0},\mathcal{C}_{0},\mathcal{T}_{t}\rangle$. We posit that the variability of the toolset is the most decisive factor for a general-purpose agent system. While memory enhances performance by recalling experience, the ability to effectively solve problems in novel domains is fundamentally limited by the availability of appropriate tools (Wang et al., 2023; Qin et al., 2024).
Figure 1: An architecture overview of Yunjue Agent.
3 Methodology
3.1 In-situ self-evolving via tool accumulation
We posit that the fundamental prerequisite for a generalist agent lies in the continuous, dynamic evolution of its toolset. To address this, we introduce an agentic workflow for in-situ self-evolving via tool accumulation.
Upon receiving a query $x_{t}$, the agent first attempts to retrieve relevant utilities from its existing repository, denoted as $\mathcal{T}_{sub}\subseteq\mathcal{T}_{t-1}$. In the absence of requisite capabilities, the agent synthesizes novel tools $\mathcal{P}_{t}$ tailored to the specific task constraints. Subsequently, empowered by the augmented toolset $\mathcal{T}_{sub}\cup\mathcal{P}_{t}$, the agent executes the task through a sequence of tool invocations. To ensure that generated tools are not merely functional but also robust and reusable, the agent engages in a self-reflection mechanism post-execution, refining the tools based on the error reports and execution traces. The process concludes with the integration of these refined tools into the global repository, updating the state to $\mathcal{T}_{t}=\mathcal{P}_{t}\cup\mathcal{T}_{t-1}$.
As the agent processes an expanding stream of queries, existing tools undergo iterative refinement while new ones are concurrently synthesized. This dual mechanism propels the system’s evolution along two dimensions: breadth (expanding functional coverage) and depth (optimizing tool robustness). Consequently, upon exposure to a sufficient volume of tasks, the tool repository is expected to reach a state of convergence, where incremental synthesis is necessitated only by outlier queries with idiosyncratic requirements.
3.2 Parallel batch evolution
While sequential query processing maximizes tool reuse and minimizes redundancy, it is inherently inefficient for large-scale benchmarks. Conversely, fully independent parallel processing fails to leverage shared knowledge accumulation across contexts. To reconcile these objectives, we propose a Parallel Batch Evolution strategy that parallelizes agent dynamics while maintaining a cohesive evolutionary trajectory.
Formally, let $\mathcal{Q}_{t}=\{q_{t,1},q_{t,2},\dots,q_{t,B}\}$ denote a batch of $B$ user queries input to the agent $\mathcal{M}_{t-1}$ at step $t$. For each query $q_{t,i}\in\mathcal{Q}_{t}$, the system synthesizes a set of local tools $\mathcal{P}_{t,i}$ that augment the global toolset $\mathcal{T}_{t-1}$ specifically for that instance. However, independent generation often yields functionally redundant tools across different queries—particularly for general-purpose utilities like web searching. This redundancy expands the tool search space, increasing the cognitive load on the agent. To mitigate this, we introduce a tool absorbing mechanism designed to cluster and consolidate utilities post-generation. Specifically, upon completion of batch $\mathcal{Q}$, the system aggregates all tools $\{\mathcal{T}_{t-1},\mathcal{P}_{t,1},\dots,\mathcal{P}_{t,B}\}$ and clusters them into disjoint groups $\{G_{j}\}$ based on functional semantic similarity. Subsequently, a merging function $\Phi$ consolidates these groups, filtering for quality and redundancy to produce a compact updated pool $\mathcal{T}_{t}=\Phi(\{G_{j}\})$. This process ensures the tool space remains streamlined, preventing retrieval ambiguity while updating the system state to $\mathcal{M}_{t}$.
Processing queries in batches offers distinct advantages. Primarily, it significantly enhances system throughput. Furthermore, analogous to mini-batch gradient descent in model optimization (Bottou, 2010; Goodfellow et al., 2016)—which mitigates gradient variance by averaging over samples—our absorbing mechanism reduces evolutionary stochasticity by merging similar tool instances. Simultaneously, this acts as a form of Best-of-N test-time scaling (Cobbe et al., 2021; Beirami et al., 2024), effectively performing multiple parallel rollouts for tool creation and selecting the optimal synthesis results for the permanent library.
3.3 The Yunjue agent system
By decoupling tool management, synthesis, and task execution, we establish a multi-agent system (Park et al., 2023) optimized for in-situ self-evolving. As illustrated in Figure 1, the architecture comprises distinct functional roles: Manager, Executor, Tool Developer, and Integrator, supported by the Aggregator and Merger for batch-level synchronization. The workflow proceeds via a collaborative mechanism designed to handle complex queries flexibly:
Upon receiving a user query $q_{t}$, the Manager orchestrates the workflow by first aligning tool capabilities with task requirements. It analyzes the system state to retrieve a relevant subset of tools $\mathcal{T}_{\text{sub}}\subseteq\mathcal{T}_{t-1}$ from the global repository. Should capability gaps arise, the Manager directs the Tool Developer to synthesize bespoke tools (implemented as Python primitives), which are immediately instantiated in the local context. Subsequently, the Executor addresses the query using the provisioned toolset, adhering to the ReAct paradigm (Yao et al., 2023). Crucially, the system supports dynamic runtime adaptation: if the Executor encounters unforeseen capability deficits during reasoning (e.g., identifying the need for a specific file parser), it suspends execution and signals the Manager. The Manager then provisions the requisite tools on-the-fly, allowing the Executor to resume seamlessly.
Figure 2: Performance comparison of Yunjue Agent against state-of-the-art agents and agentic foundation models. Our method is highlighted in cyan, and the backend model (Gemini 3 Pro) appears in orange. *Gemini 3 Pro denotes our implementation with a Python interpreter.
Upon task completion, the Integrator consolidates the execution history and intermediate outputs to formulate a comprehensive response $a_{t}$. Through this iterative process, the Yunjue agent evolves from state $\mathcal{M}_{t-1}$ to $\mathcal{M}_{t}$, effectively updating the tool pool to $\mathcal{T}_{t}$ for future evolution.
Parallel evolution with batch processing. To optimize evolutionary throughput, we implement the parallel batch evolution strategy. Given a query batch $\mathcal{Q}_{t}=\{q_{t,1},q_{t,2},\dots,q_{t,B}\}$, the system processes each instance concurrently, granting all agents shared access to the global tool repository $\mathcal{T}_{t-1}$. This parallel execution allows each query $q_{t,i}$ to independently synthesize a local toolset $\mathcal{P}_{t,i}$ tailored to its specific execution context.
Upon the completion of the batch, the tool absorbing mechanism is invoked to consolidate the dispersed local pools with the global state. Specifically, the system aggregates the union of all toolsets $\{\mathcal{T}_{t-1},\mathcal{P}_{t,1},\mathcal{P}_{t,2},\dots,\mathcal{P}_{t,B}\}$. The LLM-based Aggregator then groups utilities based on functional semantic similarity. Subsequently, another LLM-based Merger is applied to each cluster to synthesize a unified, canonical tool that encapsulates collective capabilities while eliminating redundancy. The resulting consolidated repository $\mathcal{T}_{t}$ serves as the initialized state for the subsequent queries.
4 Experiment Setup
Datasets. To demonstrate the generalizability of our approach across diverse domains and task complexities, we conduct comprehensive evaluations on four complementary benchmarks, each targeting distinct professional scenarios: (i) HLE (Humanity’s Last Exam) (Phan et al., 2025), a frontier multi-modal benchmark featuring expert-level questions across mathematics, humanities, and natural sciences, designed to assess advanced reasoning at the boundary of human knowledge; (ii) DeepSearchQA (DSQA) (Gupta et al., 2025), which challenges agents’ ability to synthesize comprehensive answers through deep web search, iterative information gathering, and multi-source evidence aggregation; (iii) xBench (Chen et al., 2025b), a Chinese profession-aligned evaluation suite where we focus on the ScienceQA (xSciQA) subset (spanning natural, social, and language sciences) and DeepSearch (xDS, 2025.10 version) subset to evaluate real-world productivity in scientific research and complex retrieval workflows, allowing us to assess cross-lingual adaptation capabilities; and (iv) FinSearchComp (FSC) (Hu et al., 2025), a bilingual (English and Chinese) benchmark targeting financial analysis through its T2 (Simple Historical Lookup) and T3 (Complex Historical Investigation) tasks, which require precise time-sensitive data retrieval and multi-step quantitative reasoning over financial documents.
Baselines. We benchmark our agent system against a comprehensive suite of proprietary and open-source systems, spanning both static and self-evolving agent paradigms. Unless otherwise noted, reported performance metrics are derived directly from original publications, technical reports, or authoritative disclosures (detailed provenance is provided in the Appendix).
5 Evaluation on open-ended evolution
5.1 Zero-start performance on cross-domain tasks
We evaluate our Yunjue Agent, initialized with an empty toolset, across all five benchmarks against state-of-the-art baselines. Results are presented in Figure 2. For datasets excluding xSciQA, foundation model baselines (e.g., GPT-5, Gemini 3 Pro) are augmented with web search and Python interpreters. Conversely, on xSciQA, where standard leaderboards prohibit external tools, we explizcitly equip our backend model with a Python interpreter to ensure a fair parity comparison.
On the heterogeneous HLE benchmark, our agent yields a significant improvement over the backend (48.0 vs. 45.8), ranking second only to GPT-5.2 Pro. Most notably, Yunjue Agent achieves state-of-the-art performance on DSQA, FSC, and xSciQA. We observe substantial absolute gains over the Gemini 3 Pro baseline, with increases of +17.4 points on DSQA (73.5 vs. 56.6) and +15.1 points on FSC (65.0 vs. 49.9), alongside a record-setting score of 76.5 on xSciQA. Finally, on xDS, our method maintains competitive performance (59.7), surpassed only by GPT-5 Pro while outperforming all other models by a significant margin. These exceptional results demonstrate our method’s superior performance across a wide spectrum of tasks, laying a solid foundation for generalized knowledge accumulation and transfer.
Complementing the performance analysis, we aggregated toolsets across all five benchmarks to examine functional utilization patterns. Figure 3 presents the invocation frequency of the top 50 tools. The distribution reveals the spontaneous emergence of high-utility fundamental functions, most notably web_search, fetch_web_content and evaluate_math_expression. The dominance of these foundational tools confirms that the system has effectively distilled generalized knowledge into versatile primitives, ensuring broad applicability across diverse task semantics.
Figure 3: Frequency distribution of the toolset evolved across five benchmarks. We report the top 50 tools, illustrating the emergence of high-generalizability primitives.
5.2 Warm start evolution in shifting domains
We further validate that the generalized knowledge accumulated by our framework is transferable to novel domains, capable of catalyzing capability evolution upon a pre-existing foundation. Specifically, we conduct sequential experiments on DSQA and xSciQA, initializing the system with the toolset derived from the HLE benchmark. HLE is selected as the foundational domain due to its substantial scale (2,500 queries) and extensive disciplinary coverage.
Table 1 presents a comparison of performance and tool synthesis across initialization strategies. We observe that performance remains consistent across domains, with a notable improvement on xSciQA, where the score increases from 76.5 to 80.2. Crucially, the number of newly synthesized tools decreases effectively by 32% on DSQA and remarkably by 100% on xSciQA. This substantial reduction demonstrates that our system’s accumulated generalized knowledge can be seamlessly transferred to disparate domains while retaining the capacity for continuous learning of novel capabilities.
Figure 4: Evolution of the tool library size relative to the cumulative number of processed queries. The experimental sequence follows the curriculum HLE $\rightarrow$ DeepSearchQA $\rightarrow$ FinSearchComp $\rightarrow$ xbench-ScienceQA $\rightarrow$ xbench-DeepSearch, highlighting the convergence of tool synthesis.
Figure 5: Venn diagram visualization of tool correspondence between zero-start and warm-start settings on DeepSearchQA. The left set comprises tools unique to the zero-start baseline ($\mathcal{T}_{\text{DSQA}}\setminus\mathcal{T}_{\text{HLE}}$), while the right set consists of incrementally generated tools in the warm-start setting ($\mathcal{T}_{\text{HLE}\rightarrow\text{DSQA}}\setminus\mathcal{T}_{\text{HLE}}$). Central intersection indicates high functional overlap. Distinct points represent individual tools, arranged by semantic similarity. Tools in the intersection share similar functionalities, while those within the central dashed lines are exact matches.
Figure 6: Evolutionary generality loss (EGL) on the HLE benchmark and accuracy on a selected dataset as a function of the number of queries. The orange curve represents the EGL trend (left axis). The blue bars show the agent’s accuracy (right axis) at different stages of the evolution (170, 1000, 1750, 2500 queries), highlighting that performance improves and stabilizes as the toolset converges.
Table 1: Performance comparison and tool synthesis statistics across different initialization strategies. ZS stands for a zero start setting, and WS denotes warm start, which leverages the HLE-evolved toolset as initialization.
Dataset Strategy Accuracy # New tools
HLE ZS 48.0 97
DeepSearchQA ZS 74 34
WS ↑0.674.6↑0.6 ↓32%23↓32%
FinSearchComp ZS 65.0 18
WS ↑0.465.4↑0.4 ↓55%8↓55%
xbench-ScienceQA ZS 76.5 13
WS ↑3.780.2↑3.7 ↓100%0↓100%
xbench-DeepSearch ZS 59.7 16
WS ↑0.960.6↑0.9 ↓100%0↓100%
Figure 4 illustrates the evolution of the tool library size relative to the cumulative number of processed queries. The overall trajectory demonstrates a global trend toward convergence. A localized surge in tool synthesis during the late phase of HLE (queries 2,000–2,400) is driven by intra-benchmark domain shifts: as query semantics transition from predominantly mathematical to social sciences and other disciplines, a targeted expansion of the toolset is necessitated. Notably, despite this semantic heterogeneity, only 97 tools were generated across the entire 2,500-HLE query corpus, evidencing the effective consolidation of acquired knowledge. Subsequently, the tool growth curves for DSQA and xSciQA exhibit a near-zero gradient. This plateau indicates that the capabilities evolved during the HLE phase possess sufficient generalization to support robust adaptation to novel domains with negligible marginal tool synthesis.
To empirically substantiate the role of tools as vehicles for knowledge crystallization and domain adaptation, we analyze the toolset intersection visualized in Figure 5. The comparative analysis reveals a substantial overlap between tools generated in the zero-start regime and those synthesized during warm-start evolution. Specifically, the recurrence of exact matches, such as read_excel, validates the system’s capacity to deterministically recover essential utilities. Furthermore, the emergence of semantically aligned variants (e.g., inspect_ods, fetch_api) highlights the agent’s adaptability in synthesizing functionally equivalent primitives tailored to specific task constraints. The predominance of these convergent tools corroborates our hypothesis of systemic robustness in shifting domains, confirming that the agent reliably gravitates toward an invariant core of capabilities to fulfill domain requirements, independent of the initialization state.
5.3 Evolutionary generality loss
To quantify the dynamics of knowledge acquisition, we propose a novel metric termed Evolutionary generality loss (EGL). Functioning analogously to the objective function in stochastic optimization, EGL serves as a real-time indicator of evolutionary convergence. Formally, EGL is defined as:
$\text{EGL}=\frac{\sum_{i=1}^{N}c_{i}}{\sum_{i=1}^{N}u_{i}}\times 1000$ (1)
where $c_{i}$ and $u_{i}$ denote the number of newly synthesized tools and tool invocations for the $i$-th query, respectively, and $N$ represents the total number of processed queries. Intuitively, the EGL value is elevated during the nascent stages of evolution, driven by the necessity of ab initio tool creation. As the system matures, the dominance of tool reuse over synthesis causes the metric to decay. Conversely, a sustained high EGL value implies that the system is failing to generalize, necessitating continuous creation. Conceptually, this metric is inversely correlated with tool generality: as the versatility of the repository increases, the EGL value decreases. To validate its efficacy, we monitored the EGL trajectory on the HLE benchmark as a function of query volume. The resulting trend in Figure 6 exhibits a distinct convergence pattern, achieving asymptotic stability after processing approximately 1,000 queries (40% of the stream).
Figure 7: Evolution dynamics under different batch sizes. (a) The number of created tools with the number of queries. (b) Average token consumption per tool use. (c) Cumulative number of tool invocations.
To empirically corroborate this convergence, we extracted system snapshots at 10%, 40%, 70%, and 100% of the evolutionary timeline and evaluated performance on a stratified random sample of 200 queries (100 each from the HLE and DSQA datasets), with further tool synthesis disabled. As illustrated in the bar chart of Figure 6, a substantial performance inflection occurs between the 10% and 40% checkpoints, signifying a rapid expansion of the system’s capability boundary. Conversely, marginal gains diminish significantly in the 40%–100% interval. This saturation provides strong evidence that the evolutionary process successfully converges, reaching a state of high generality where incremental tool synthesis yields diminishing returns.
Table 2: Experimental results on a subset of 200 selected queries under varying batch sizes. “Tool succ. rt.” denotes the tool execution success rate, while “Avg #tokens for tools” quantifies the average token consumption per tool invocation. The condition $\infty$” represents that the backend model operates exclusively with a Python interpreter, processing queries independently without tool accumulation.
Batch size Accuracy Tool succ. rt. Avg #tokens for tools
2 48.0 99.6% 109.9
4 50.0 99.9% 130.3
8 47.5 99.6% 188.4
16 51.5 99.1% 108.4
$\infty$ 40.0 81.8% 518.2
5.4 Ablation study
To elucidate the factors underpinning the efficacy of tool evolution and to ablate key system parameters, we conduct comparative experiments using the identical subset of 200 queries employed in the EGL evaluation.
Why does tool evolution work? To validate the fundamental utility of tool evolution, we establish a baseline comparison against an agent sharing an identical backend (GPT-5) but equipped solely with a Python interpreter. While the interpreter allows for ad-hoc code execution and iterative debugging based on error feedback, it fundamentally lacks the capacity for knowledge crystallization and reuse.
We evaluate our agent under a zero-start regime with varying batch sizes, contrasting it against the Python interpreter baseline (denoted as $B\to\infty$). The condition $\infty$” represents the theoretical asymptotic limit of infinite batch size, where the backend model operates exclusively with a Python interpreter. In this setting, queries are processed independently, precluding any tool accumulation, clustering, or shared knowledge transfer. As detailed in Table 2, the Python-only baseline yields a markedly inferior score of 40.0, compared to the 47.5–51.5 range achieved by our method. To elucidate the underlying causes, we analyzed the tool success rate (successful invocations per total attempts) and token efficiency (average tokens per invocation). The baseline demonstrates deficiencies in both metrics, exhibiting a success rate of only 81.8% while consuming an average of 518.2 tokens per invocation. In stark contrast, our system consistently maintains success rates exceeding 99% with significantly reduced token overhead. This divergence underscores that evolved tools are not only more robust but also more efficient in information processing. We posit that the high execution error rate and the verbosity of raw code generation induce severe contextual contamination, thereby degrading the reasoning capabilities of the Python baseline.
Table 3: Performance comparison and number of created tools of different backend models on DeepSearchQA and FinSearchComp benchmarks.
Backend DeepSearchQA FinSearchComp
Accuracy # Tools Accuracy # Tools
Gemini 3 Pro 74.0 16 65.0 18
GPT-5 70.6 75 68.2 25
GPT-5-mini 42.5 47 45.5 13
How does batch size impact parallel evolution? We further investigate the role of batch size $B$ in mediating the trade-off between evolutionary velocity and convergence stability. As illustrated in Figure 7(a), larger batch configurations induce a steeper initial gradient in tool synthesis. This suggests that high-concurrency processing facilitates the rapid identification of common patterns and the accelerated crystallization of foundational primitives. Crucially, despite varied initial velocities, the terminal library sizes remain comparable across all settings, confirming that the system invariably converges to a consistent equilibrium of capabilities regardless of batch magnitude.
Figure 7(b) tracks the temporal evolution of computational overhead. Even in the nascent stages, our parallelized architecture yields reduced token consumption relative to the baseline (w/ py exec). As the system matures, the average token cost per invocation demonstrates a precipitous decay, eventually stabilizing at an asymptotic minimum. Note that localized fluctuations in this downward trend correspond to the sporadic synthesis of novel tools required for outlier queries. This overall convergence signifies the system’s successful phase transition from expensive ab initio code generation to efficient, low-latency tool reuse.
We observe in Figure 7(b) that the average token consumption per tool use after convergence for a batch size of 8 is slightly higher than other settings. This is because, as shown in Figure 7(a), it generates the largest number of tools, and Figure 7(c) indicates it has the highest cumulative tool invocations. We attribute this to the stochastic nature of LLM execution, which leads to varying agent behaviors under different settings. Future work will investigate strategies to align the convergence states of the system across different batch sizes.
How does backend model selection affect evolution? We evaluate the influence of the underlying foundation model on system performance across the HLE and DSQA datasets, as detailed in Table 3. Results indicate that when instantiated with GPT-5, our agent surpasses the majority of established baselines. Notably, the lightweight GPT-5-mini variant achieves competitive performance, demonstrating significant utility despite its reduced parameter count. Crucially, we employ a unified prompt strategy across all configurations without architecture-specific tuning. These findings underscore the model agnosticism and generalization capability of our framework, confirming its robustness across diverse backend scales.
6 Related Work
Self-evolving agents. To transcend the limitations of static designs, general-purpose agent systems must adapt to open-ended, interactive environments in real-time (Gao et al., 2025). Self-evolving agents achieve this by autonomously refining internal components, categorized into workflow optimization, model parameter updates, context management, and tool synthesis. Specifically, MAS-GPT (Ye et al., 2025) and AFlow (Zhang et al., 2025a) generate query-specific workflows, while TT-SI (Acikgoz et al., 2025) and SCA (Zhou et al., 2025) update parameters by generating challenging training problems. Furthermore, ELL (Cai et al., 2025) and Zhang et al. (2025b) evolve by accumulating environmental interaction experience as context. Regarding tool synthesis, LIVE-SWE-agent (Xia et al., 2025) and STELLA (Jin et al., 2025) enable on-the-fly tool creation for software engineering and biomedical research, respectively. Nevertheless, most self-evolving methods necessitate explicit training. Crucially, even approaches closest to our work—those autonomously creating tools at test-time—remain confined to specific domains and lack tool reuse mechanisms, hindering their general applicability.
Tool evolution agents. The paradigm of autonomous tool synthesis has garnered significant attention in recent literature. Pioneering systems such as Voyager (Wang et al., 2023), STELLA (Jin et al., 2025), and LIVE-SWE-agent (Xia et al., 2025) empower agents to generate executable tools for embodied control or software engineering tasks. Notably, analyses of the tool embedding space in LIVE-SWE-agent (Xia et al., 2025) reveal distinct clustering patterns among functional equivalents; this observation directly informed the design of our absorbing mechanism, which utilizes clustering to consolidate redundant utilities. Regarding tool reuse, Alita (Qiu et al., 2025) facilitates the persistence of successful tools, though its optimization trajectory is predicated on ground-truth supervision. In the context of inference-time adaptation, Lu et al. (2026) proposed a test-time tool evolution framework governed by a cost-sensitive objective function, yet its application remains narrowly scoped to scientific reasoning. Collectively, these approaches prioritize the optimization of existing assets or domain-specific synthesis, distinguishing them from our framework’s focus on self-evolution in open-ended environments.
General agent systems. Recent research has sought to construct general-purpose agents capable of cross-domain operation. DeepAgent (Li et al., 2025) utilizes a scalable toolset to address diverse tasks but lacks the capacity for autonomous environmental adaptation, leading to performance limitations. Agent KB (Tang et al., 2025) attempts to solve cross-domain problems by summarizing execution experiences from other agents; however, its performance is constrained by a deficiency in necessary tools. In contrast, our approach achieves generality by continuously creating, reusing, and optimizing tools in-situ, effectively overcoming the limitations of static toolsets and experience-based adaptation.
7 Discussion and future work
While this study establishes the efficacy of in-situ self-evolution through tool synthesis, several avenues for future research merit rigorous inquiry to fully realize the potential of autonomous agents.
Paradigm parallels: towards system-level pre-training for agentic systems. The seamless transition and performance consistency observed between our zero-start and warm-start settings offer more than just empirical validation; they suggest a fundamental paradigm shift. The clear convergence curves of the tool library indicate that “task-solving capability” is not merely a collection of ad-hoc heuristics, but a generalizable pattern that can be learned and distilled.
This implies that the field is approaching a “pre-training and post-training” era for agentic systems, mirroring the trajectory of LLMs. We envision a future where multi-agent systems undergo system-level pre-training on massive, broad-spectrum task datasets. This process would allow agents to distill a “foundation toolset”—a set of converged, generalizable primitives—before deployment. Consequently, such pre-trained agents would possess intrinsic generalization capabilities, enabling them to tackle novel downstream tasks largely through the composition of existing reliable tools, thereby minimizing or even eliminating the need for expensive test-time evolution. Formalizing the methodologies for this “agentic pre-training” is a critical priority for future research.
Co-evolution of memory and workflow. Currently, our framework validates the self-evolving paradigm primarily through tool generation. However, tools constitute a necessary but insufficient condition for scenarios demanding high personalization or complex process management. For instance, personalized assistants require persistent, evolving memory structures to align with user preferences, while intricate tasks (e.g., deep research) necessitate structured workflow evolution—such as generating bespoke planning protocols for multi-agent coordination. A critical direction is extending the evolutionary mechanism to encompass the co-evolution of memory architectures and workflow policies, enabling the system to adapt its internal state and execution logic alongside its functional capabilities.
Evolutionary stability and regularization. The stability of tool evolution warrants further investigation. Due to the inherent stochasticity of LLM generation, toolsets can exhibit variance across experimental runs, as evidenced by our batch size ablation. Ensuring the consistent convergence of the tool library is vital for system reliability. Future work will focus on developing regularization strategies to guarantee the determinism of the evolutionary process in open-ended environments.
Optimization of parallel batch evolution. Finally, the nuances of the batch evolution strategy present rich opportunities for optimization:
  • Curriculum learning effects: As shown in Figure 4, the sequence of incoming queries significantly influences the convergence trajectory; delaying queries that require foundational primitives can impede system maturation. Investigating optimal query ordering is a key direction.
  • Intra-batch diversity trade-offs: Intra-batch diversity presents a complex dynamic. Low diversity allows the absorbing mechanism to function as a form of Best-of-N test-time scaling, leveraging redundancy to select the optimal implementation—a contrast to traditional gradient training where high batch diversity is preferred. Balancing this quality-assurance benefit against the efficiency of capability evolution is crucial.
  • Adaptive scheduling: Dynamic batch sizing offers a promising avenue for optimization. Smaller batch sizes in early stages could foster the robust consolidation of general knowledge via redundancy, while larger batches in the post-convergence phase could maximize throughput for corner cases. Developing autonomous agents capable of dynamic batch scheduling based on convergence signals remains a worthy research direction.
8 Conclusion
In this work, we presented the In-Situ Self-Evolving framework, enabling LLM-based agents to autonomously adapt and evolve within open-ended environments. Through the Yunjue Agent and the proposed Parallel Batch Evolution strategy, we demonstrated that treating tools as dynamic vehicles for knowledge crystallization allows for robust zero-start learning and the emergence of generalized, transferable capabilities. Our empirical results validate that this approach not only achieves state-of-the-art performance across heterogeneous benchmarks but also significantly outperforms static baselines in both efficiency and adaptability. Furthermore, warm-start experiments confirm that the accumulated generalized knowledge can be seamlessly transferred to novel domains, enabling the agent to adapt to shifting environments with minimal marginal tool synthesis. We release our codebase, evaluation and system traces to the open-source community to facilitate further research. Broadly, future work will aim to unify tool evolution with memory and workflow adaptation, advancing toward more autonomous and general-purpose agentic systems.
Appendix A Experiment setup
Table 4: Statistics of the datasets.
Dataset Domain Language # Queries
HLE General reasoning EN 2,500
DeepSearchQA General QA EN 900
FinSearchComp (T1 & T2) Financial search EN & CH 391
xBench-ScienceQA General science CH 100
xBench-DeepSearch General research CH 100
Figure 8: EGL dynamics on the DSQA and FSC datasets.
Data statistics are summarized in Table 4.
Evaluation protocols. We follow the standard evaluation protocols released with each dataset:
Baselines sources. We collect baseline results from their official reports or leaderboards. For HLE, results for Gemini Deep Research and Gemini 3 Pro are from (Google, 2025); MiroThinker-v1.0-72B and Tongyi DeepResearch are from (Team et al., 2025); GPT-5.2 Pro is from (OpenAI, 2025); Seed-1.8 is from (ByteDance, 2025); and Claude 4.5 Opus is from (Anthropic, 2025). DeepSearchQA results are from (Gupta et al., 2025). FinSearchComp results are from (ByteDance, 2025). For xBench, results on ScienceQA and DeepSearch are retrieved from (xBench Team, 2025b) and (xBench Team, 2025a), respectively.
Implementation details. To ensure experimental consistency, all nodes in our system are instantiated using the same backend LLM: Gemini 3 Pro. The batch size is set to $B=16$ for all datasets. The Tool Developer module is explicitly powered by Codex (Zaremba et al., 2021) to facilitate robust code generation. We main a fixed temperature of 0.7 across all LLM invocations. To simplify the use of multimodal capabilities, we encapsulate image processing functionality into a dedicated tool, ensuring utilization of the same backend LLM. Our agent system is built based on LangGraph (LangChain, 2024). All prompts in our system adopt the Markdown template format supported by Jinja (The Pallets Projects, 2007).
Appendix B More evaluation results
More results on evolutionary generality loss (EGL). Figure 8 tracks the EGL trajectories across the DSQA and FSC benchmarks, while Figure 9 dissects the impact of batch size on optimization behavior. Collectively, these visualizations validate the asymptotic stability of our approach across varying experimental conditions.
Figure 9: EGL dynamics under different batch sizes.
Figure 9 demonstrates the EGL dynamics under varying batch sizes, indicating the robustness and stability of our training process.
Technical insights. We observed that Gemini 3 Pro possesses strong reasoning capabilities and internal knowledge, often exhibiting a sense of “confidence”, which is manifested by its tendency to rely on fewer tools to complete tasks. However, it frequently suffers from hallucinations and often exhibits weaker instruction-following capabilities, necessitating iterative prompt engineering. In contrast, GPT-5 and GPT-5-mini often appear more “cautious”, typically requiring the invocation of more tools, planning, and reasoning steps to iteratively verify task results. This leads to increased tool creation and longer execution histories. This phenomenon is particularly pronounced in GPT-5 (as shown in Table 3, the GPT-5 series clearly creates more tools). Nevertheless, the GPT-5 series demonstrates exceptional instruction-following abilities; adding constraints to their prompts generally yields behavior consistent with expectations.
Appendix C Case study
Figure 10: Yunjue agent execution pipeline. Upon receiving a query, the agent autonomously synthesizes, deploys, and iteratively refines tools to derive the final response, seamlessly integrating generation and execution.
Figure 11: Illustration of tool absorbing mechanism. Following batch execution, functionally analogous tools are identified via clustering and consolidated into a generalized, compact toolset, effectively pruning redundancy.
Figure 12: Python-only baseline failure mode. Unlike our approach, the baseline indiscriminately accumulates raw execution traces—including erroneous attempts—within the context window, leading to severe contextual contamination and reasoning degradation.
Appendix D Prompts
Manager Tool Developer Executor Integrator Aggregator Merger