### v2

# Yunjue Agent Tech Report: A Fully Reproducible, Zero-Start In-Situ Self-Evolving Agent System for Open-Ended Tasks

Abstract
Conventional agent systems often struggle in open-ended environments where task distributions continuously drift and external supervision is scarce. Their reliance on static toolsets or offline training renders capability boundaries rigid, preventing adaptation to evolving dynamics. To address this, we propose the In-situ Self-Evolving paradigm. This approach treats sequential task interactions as a continuous stream of experience, enabling the system to distill short-term execution feedback into long-term, reusable capabilities without ground-truth labels. We identify tool evolution as the critical pathway for this capability expansion. Accordingly, we develop the Yunjue Agent, a system that iteratively synthesizes, optimizes, and reuses tools to navigate emerging challenges. To optimize evolutionary efficiency, we introduce a Parallel Batch Evolution strategy. Furthermore, we propose a novel metric to monitor evolutionary convergence, serving as a function analogous to training loss in conventional optimization. Empirical evaluations across five diverse benchmarks under a zero-start setting demonstrate significant performance gains over proprietary baselines and validate the effectiveness of our proposed convergence metric. Additionally, complementary warm-start evaluations confirm that the accumulated general knowledge can be seamlessly transferred to novel domains. We open-source our codebase, system traces, and evolved tools to facilitate future research in resilient, self-evolving intelligence.
Machine Learning, ICML
1 Introduction
The rapid advancement of Large Language Models (LLMs) has catalyzed a paradigm shift in the pursuit of Artificial General Intelligence (AGI) (Guo et al., 2024), fundamentally expanding the AI frontier across diverse daily and scientific domains. While these developments have birthed versatile agents, a critical limitation persists: most contemporary systems remain confined within predefined operational envelopes. This reliance on static functional boundaries renders agents inherently fragile when confronted with the stochasticity of open-world environments. We posit that true AGI demands not merely broad-domain competence, but the capacity for self-adaptive evolution—dynamically reshaping capabilities to surmount novel challenges within open-ended contexts. Consequently, architecting a system capable of both cross-domain execution and recursive self-evolution has emerged as a paramount challenge in the quest for resilient intelligence.
Within the framework of LLM-based agents, such adaptive capability typically rests on three pillars: workflow, context, and tool. Recently, the field has witnessed a surge in self-evolving agents (Tao et al., 2024; Gao et al., 2025), designed to autonomously enhance capabilities without human intervention, achieving progress in workflow adaptation (Zhang et al., 2025b), context management (Tang et al., 2025), and tool utilization (Wang et al., 2023). However, existing paradigms often rely on offline supervision (Acikgoz et al., 2025) or remain confined to narrow domains (Wang et al., 2023; Jin et al., 2025). In supervision-scarce, open-ended environments, a more autonomous paradigm is imperative. We argue that among the three pillars of agent systems, while workflows refine expertise and context aligns preference, the toolset constitutes the fundamental prerequisite for novel environments. Without requisite functional tools, even sophisticated planning or memory becomes computationally paralyzed. Furthermore, tool functionality is verifiable via rigorous execution: a program either succeeds or throws an exception. This robust binary signal enables autonomous self-correction even absent human intervention, positioning tools as the optimal starting point for in-situ evolution.
To address the open-ended challenges, we propose the In-situ Self-Evolving (ISE) framework. This paradigm treats task interactions as a continuous stream for accumulating generalized knowledge, initiating with a zero-start phase. We develop the Yunjue Agent, utilizing a multi-agent architecture (Manager, Tool Developer, Executor, etc.) (Qian et al., 2024) to synthesize bespoke Python primitives on-the-fly when existing capabilities are insufficient. To further optimize efficiency, we introduce a Parallel Batch Evolution strategy, which distills high-quality tools from concurrent executions to accelerate convergence. Furthermore, we propose a novel metric to monitor evolutionary convergence, functioning analogously to training loss. We open-source our codebase, complete interaction traces, and the library of evolved tools to provide a transparent foundation for future research in resilient, self-evolving intelligence.
Comprehensive empirical evaluations across five benchmarks validate the Yunjue Agent’s efficacy. Results demonstrate: (i) State-of-the-art performance: In zero-start settings, autonomous tool evolution outperforms proprietary baselines; (ii) Cross-domain transferability: Additional warm-start evaluations show that accumulated general knowledge adapts effectively to novel domains; (iii) Robust convergence monitoring: Our proposed metric effectively serves as a stability proxy for evolution; and (iv) Cost-effectiveness: The lightweight design ensures economical operation.
2 In-situ self-evolving agents
An agent system can be formally defined as a tuple $\mathcal{M}=\langle\mathcal{W},\mathcal{C},\mathcal{T}\rangle$ of workflow, context and tools (Gao et al., 2025). In this system, the workflow $\mathcal{W}$ is typically structured as a directed graph $\mathcal{G}=(V,E)$, where nodes $V$ represent LLM-based agents and edges $E$ denote the flow of information. $\mathcal{C}$ is the set of contexts, where each context element can be a prompt template or a memory buffer encompassing dynamic conversation history. Finally, $\mathcal{T}$ denotes the set of tools available for the agents to execute specific tasks.
To enable agents to adapt to new environments like humans, recent work focuses on self-evolving methods (Tao et al., 2024; Gao et al., 2025). However, these methods typically require an iterative training process to update certain components (He et al., 2025; Chen et al., 2025a) or are confined in specific areas (Xia et al., 2025; Jin et al., 2025), such as generating query-specific workflows (Zhang et al., 2025b; Ye et al., 2025), refining context $\mathcal{C}$ (e.g., prompt and memory) (Zhang et al., 2025b; Tang et al., 2025) or adjusting toolset (Wang et al., 2024, 2023). In contrast, we aim to explore whether an agent system can adapt to the environment by continuously updating its components in-situ, where there is no access to external supervision signals. This motivation leads to a new paradigm of agentic evolution.
In-situ self-evolving of an agent system. Given a sequence of queries $\{x_{1},x_{2},\dots,x_{T}\}$, the agent evolves dynamically from $\mathcal{M}_{0}$ to $\mathcal{M}_{T}$ as it processes each. Specifically, after completing the $t$-th query $x_{t}$, the system updates its configuration for the next query, which can be formalized as a transition from $\mathcal{M}_{t-1}=\langle\mathcal{W}_{t-1},\mathcal{C}_{t-1},\mathcal{T}_{t-1}\rangle$ to $\mathcal{M}_{t}=\langle\mathcal{W}_{t},\mathcal{C}_{t},\mathcal{T}_{t}\rangle$, where any component of the tuple—the workflow structure, context, or toolset—may be modified based on internal feedback or experience gained from the previous interaction. Unlike self-evolving agents that maximize a specific objective function through a series of training steps (Gao et al., 2025), in-situ self evolving operates during the inference phase where ground truth information is unavailable.
To ensure a focused scope, we fix the workflow $\mathcal{W}$. Furthermore, given that Agent KB (Tang et al., 2025) has attempted to optimize memory through cross-domain experience to enhance agent performance, we also treat the context $\mathcal{C}$ as a fixed component [^1] Although the execution flow can be dynamic during inference due to branch control and model inputs are populated at runtime, the underlying workflow structure and prompt templates are pre-defined. Thus, we consider these configurations fixed., reducing the evolving system state to $\mathcal{M}_{t}=\langle\mathcal{W}_{0},\mathcal{C}_{0},\mathcal{T}_{t}\rangle$. We posit that the variability of the toolset is the most decisive factor for a general-purpose agent system. While memory enhances performance by recalling experience, the ability to effectively solve problems in novel domains is fundamentally limited by the availability of appropriate tools (Wang et al., 2023; Qin et al., 2024).
Figure 1: An architecture overview of Yunjue Agent.
3 Methodology
3.1 In-situ self-evolving via tool accumulation
We posit that the fundamental prerequisite for a generalist agent lies in the continuous, dynamic evolution of its toolset. To address this, we introduce an agentic workflow for in-situ self-evolving via tool accumulation.
Upon receiving a query $x_{t}$, the agent first attempts to retrieve relevant utilities from its existing repository, denoted as $\mathcal{T}_{sub}\subseteq\mathcal{T}_{t-1}$. In the absence of requisite capabilities, the agent synthesizes novel tools $\mathcal{P}_{t}$ tailored to the specific task constraints. Subsequently, empowered by the augmented toolset $\mathcal{T}_{sub}\cup\mathcal{P}_{t}$, the agent executes the task through a sequence of tool invocations. To ensure that generated tools are not merely functional but also robust and reusable, the agent engages in a self-reflection mechanism post-execution, refining the tools based on the error reports and execution traces. The process concludes with the integration of these newly synthesized tools into the global repository, updating the state to $\mathcal{T}_{t}=\mathcal{P}_{t}\cup\mathcal{T}_{t-1}$.
As the agent processes an expanding stream of queries, existing tools undergo iterative refinement while new ones are concurrently synthesized. This dual mechanism propels the system’s evolution along two dimensions: breadth (expanding functional coverage) and depth (optimizing tool robustness). Consequently, upon exposure to a sufficient volume of tasks, the tool repository is expected to reach a state of convergence, where incremental synthesis is necessitated only by outlier queries with idiosyncratic requirements.
3.2 Parallel batch evolution
To balance processing efficiency with shared knowledge accumulation, we propose a Parallel Batch Evolution strategy. While sequential processing maximizes reuse, it lacks scalability; conversely, independent parallelization sacrifices shared learning. Our approach reconciles these by parallelizing execution while maintaining a cohesive evolutionary trajectory.
Formally, let $\mathcal{Q}_{t}=\{q_{t,1},q_{t,2},\dots,q_{t,B}\}$ denote a batch of $B$ user queries input to the agent $\mathcal{M}_{t-1}$ at step $t$. For each query $q_{t,i}\in\mathcal{Q}_{t}$, the system synthesizes a set of local tools $\mathcal{P}_{t,i}$ that augments the global toolset $\mathcal{T}_{t-1}$ specifically for that instance. However, independent generation often yields functionally redundant tools across different queries—particularly for general-purpose utilities like web searching. This redundancy expands the tool search space, increasing the cognitive load on the agent. To mitigate this, we introduce a tool absorbing mechanism designed to cluster and consolidate utilities post-generation. Specifically, upon completion of batch $\mathcal{Q}_{t}$, the system aggregates all tools $\{\mathcal{T}_{t-1},\mathcal{P}_{t,1},\dots,\mathcal{P}_{t,B}\}$ and clusters them into disjoint groups $\{G_{j}\}$ based on functional semantic similarity. Subsequently, a merging function $\Phi$ consolidates these groups, filtering for quality and redundancy to produce a compact updated pool $\mathcal{T}_{t}=\Phi(\{G_{j}\})$. This process ensures the tool space remains streamlined, preventing retrieval ambiguity while updating the system state to $\mathcal{M}_{t}$.
Processing queries in batches offers distinct advantages. Primarily, it significantly enhances system throughput. Furthermore, analogous to mini-batch gradient descent in model optimization (Bottou, 2010; Goodfellow et al., 2016)—which mitigates gradient variance by averaging over samples—our absorbing mechanism reduces evolutionary stochasticity by merging similar tool instances. Simultaneously, this acts as a form of Best-of-N test-time scaling (Cobbe et al., 2021; Beirami et al., 2024), effectively performing multiple parallel rollouts for tool creation and selecting the optimal synthesis results for the permanent library.
3.3 The Yunjue Agent system
By decoupling tool management, synthesis, and task execution, we establish a multi-agent system (Park et al., 2023) optimized for in-situ self-evolving. As illustrated in Figure 1, the architecture comprises distinct functional roles: Manager, Executor, Tool Developer, and Integrator, supported by the Aggregator and Merger for batch-level synchronization. The workflow proceeds via a collaborative mechanism designed to handle complex queries flexibly:
Upon receiving a user query $q_{t}$, the Manager orchestrates the workflow by first aligning tool capabilities with task requirements. It analyzes the system state to retrieve a relevant subset of tools $\mathcal{T}_{\text{sub}}\subseteq\mathcal{T}_{t-1}$ from the global repository. Should capability gaps arise, the Manager directs the Tool Developer to synthesize bespoke tools (implemented as Python primitives), which are immediately instantiated in the local context. Subsequently, the Executor addresses the query using the provisioned toolset, adhering to the ReAct paradigm (Yao et al., 2023). Crucially, the system supports dynamic runtime adaptation: if the Executor encounters unforeseen capability deficits during reasoning (e.g., identifying the need for a specific file parser), it suspends execution and signals the Manager. The Manager then provisions the requisite tools on-the-fly, allowing the Executor to resume seamlessly.
Figure 2: Performance comparison of Yunjue Agent against state-of-the-art agents and agentic foundation models. Our method is highlighted in cyan, and the backend model (Gemini 3 Pro) appears in orange. *Gemini 3 Pro denotes our implementation with a Python interpreter.
Upon task completion, the Integrator consolidates the execution history and intermediate outputs to formulate a comprehensive response $a_{t}$. Through this iterative process, the Yunjue Agent evolves from state $\mathcal{M}_{t-1}$ to $\mathcal{M}_{t}$, effectively updating the tool pool to $\mathcal{T}_{t}$ for future evolution.
Parallel evolution with batch processing. To optimize evolutionary throughput, we implement the parallel batch evolution strategy. Given a query batch $\mathcal{Q}_{t}=\{q_{t,1},q_{t,2},\dots,q_{t,B}\}$, the system processes each instance concurrently, granting all agents shared access to the global tool repository $\mathcal{T}_{t-1}$. This parallel execution allows each query $q_{t,i}$ to independently synthesize a local toolset $\mathcal{P}_{t,i}$ tailored to its specific execution context.
Upon the completion of the batch, the tool absorbing mechanism is invoked to consolidate the dispersed local pools with the global state. Specifically, the system aggregates the union of all toolsets $\{\mathcal{T}_{t-1},\mathcal{P}_{t,1},\mathcal{P}_{t,2},\dots,\mathcal{P}_{t,B}\}$. The LLM-based Aggregator then groups utilities based on functional semantic similarity. Subsequently, another LLM-based Merger is applied to each cluster to synthesize a unified, canonical tool that encapsulates collective capabilities while eliminating redundancy. The resulting consolidated repository $\mathcal{T}_{t}$ serves as the initialized state for the subsequent queries. The pseudo-code of the overall process is provided in Appendix B.
4 Experiment setup
Datasets. To demonstrate the generalizability of our approach across diverse domains and task complexities, we conduct comprehensive evaluations on five complementary benchmarks, each targeting distinct professional scenarios: (i) HLE (Humanity’s Last Exam) (Phan et al., 2025), a frontier multi-modal benchmark featuring expert-level questions across mathematics, humanities, and natural sciences, designed to assess advanced reasoning at the boundary of human knowledge; (ii) DeepSearchQA (DSQA) (Gupta et al., 2025), which challenges agents’ ability to synthesize comprehensive answers through deep web search, iterative information gathering, and multi-source evidence aggregation; (iii) xBench (Chen et al., 2025b), a Chinese profession-aligned evaluation suite where we focus on the ScienceQA (xSciQA) subset (spanning natural, social, and language sciences) and DeepSearch (xDS, 2025.10 version) subset to evaluate real-world productivity in scientific research and complex retrieval workflows, allowing us to assess cross-lingual adaptation capabilities; and (iv) FinSearchComp (FSC) (Hu et al., 2025), a bilingual (English and Chinese) benchmark targeting financial analysis through its T2 (Simple Historical Lookup) and T3 (Complex Historical Investigation) tasks, which require precise time-sensitive data retrieval and multi-step quantitative reasoning over financial documents.
Baselines. We benchmark our agent system against a comprehensive suite of proprietary and open-source systems, spanning both static and self-evolving agent paradigms. Unless otherwise noted, reported performance metrics are derived directly from original publications, technical reports, or authoritative disclosures (detailed provenance is provided in the Appendix C).
Figure 3: Invocation frequency distribution of the toolset evolved across five benchmarks. We report the top 50 tools, illustrating the emergence of high-generalizability primitives.
5 Evaluation on open-ended evolution
5.1 Zero-start performance on cross-domain tasks
We evaluate our Yunjue Agent, initialized with an empty toolset, across all five benchmarks against state-of-the-art baselines. Results are presented in Figure 2. For datasets excluding xSciQA, foundation model baselines (e.g., GPT-5, Gemini 3 Pro) are augmented with web search and Python interpreters. Conversely, on xSciQA, where standard leaderboards prohibit external tools, we explicitly equip our backend model with a Python interpreter to ensure a fair parity comparison. We provide more detailed results in Appendix D, including the standard deviation of accuracy, token consumption and cost.
On the heterogeneous HLE benchmark, our agent yields a significant improvement over the backend (48.0 vs. 45.8), ranking second only to GPT-5.2 Pro. Most notably, Yunjue Agent achieves state-of-the-art performance on DSQA, FSC, and xSciQA. We observe substantial absolute gains over the Gemini 3 Pro baseline, with increases of +16.9 points on DSQA (73.5 vs. 56.6) and +15.1 points on FSC (65.0 vs. 49.9), alongside a record-setting score of 76.5 on xSciQA. Finally, on xDS, our method maintains competitive performance (59.7), surpassed only by GPT-5 Pro while outperforming all other models by a significant margin. These exceptional results demonstrate our method’s superior performance across a wide spectrum of tasks, laying a solid foundation for generalized knowledge accumulation and transfer.
Complementing the performance analysis, we aggregated toolsets across all five benchmarks to examine functional utilization patterns. Figure 3 presents the invocation frequency of the top 50 tools. The distribution reveals the spontaneous emergence of high-utility fundamental functions, most notably web_search, fetch_text_url and evaluate_expression_math. The dominance of these foundational tools confirms that the system has effectively distilled generalized knowledge into versatile primitives, ensuring broad applicability across diverse task semantics.
Although the Yunjue Agent integrates modules such as the Manager and Tool Developer around the Executor, the Executor acts as the primary entity for execution and token consumption. In Appendix D, we show that across five datasets, the Executor accounts for approximately 95% of the total token consumption. Additionally, since tools do not need to be repeatedly created, a significant amount of token consumption is also saved. Consequently, our system’s overall overhead is comparable to that of a standard ReAct agent equipped with a Python interpreter, a finding further corroborated in Table 2.
5.2 Warm-start evolution in shifting domains
We validate the transferability of accumulated knowledge to novel domains by initializing the system on DSQA, all the way to FSC, xSciQA and xDS, using the toolset derived from the HLE benchmark. HLE is selected as the foundational domain due to its substantial scale (2,500 queries) and extensive disciplinary coverage. Table 1 shows consistent performance across domains, with xSciQA scores improving from 76.5 to 80.2. Crucially, new tool synthesis drops by 55% on FSC and 100% on xSciQA and xDS, demonstrating effective transfer with minimal marginal cost.
Table 1: Performance comparison and tool synthesis statistics across different initialization strategies. ZS stands for a zero start setting, and WS denotes warm start, which leverages the HLE-evolved toolset as initialization.
Dataset Strategy Accuracy # New tools
HLE ZS 48.0 97
DeepSearchQA ZS 73.5 34
WS ↑1.174.6↑1.1 ↓32%23↓32%
FinSearchComp ZS 65.0 18
WS ↑0.465.4↑0.4 ↓55%8↓55%
xbench-ScienceQA ZS 76.5 13
WS ↑3.780.2↑3.7 ↓100%0↓100%
xbench-DeepSearch ZS 59.7 16
WS ↑0.960.6↑0.9 ↓100%0↓100%
Figure 4: Venn diagram comparing unique zero-start tools ($\mathcal{T}_{\text{DSQA}}\setminus\mathcal{T}_{\text{HLE}}$, left) and incremental warm-start tools ($\mathcal{T}_{\text{HLE}\rightarrow\text{DSQA}}\setminus\mathcal{T}_{\text{HLE}}$, right) on DeepSearchQA. Points represent semantically clustered tools. The intersection (71%) denotes functional overlap, with dashed lines marking exact matches, while the difference set accounts for 29%.
Figure 5 illustrates the evolution of the tool library size relative to the cumulative number of processed queries. The overall trajectory demonstrates a global trend toward convergence. A localized surge in tool synthesis during the late phase of HLE (queries 2,000–2,400) is driven by intra-benchmark domain shifts: as query semantics transition from predominantly mathematical to social sciences and other disciplines, a targeted expansion of the toolset is necessitated. Notably, despite this semantic heterogeneity, only 97 tools were generated across the entire 2,500-HLE query corpus, evidencing the effective consolidation of acquired knowledge. Subsequently, tool growth trajectories across secondary datasets exhibit near-zero slopes, indicating that HLE-evolved capabilities facilitate robust adaptation to novel domains with negligible marginal synthesis.
To substantiate tools as vehicles for knowledge crystallization and domain adaptation, Figure 4 visualizes the intersection between zero-start and warm-start toolsets. The recurrence of exact matches (e.g., read_excel) validates the capacity to deterministically recover essential utilities. Furthermore, semantically aligned variants (e.g., seed_http_request vs. fetch_api_data, see Appendix E) highlight adaptive functional equivalence. This convergence confirms the agent gravitates toward an invariant core of capabilities independent of initialization.
Figure 5: Evolution of the tool library size relative to the cumulative number of processed queries. The experimental sequence follows the curriculum HLE $\rightarrow$ DeepSearchQA $\rightarrow$ FinSearchComp $\rightarrow$ xbench-ScienceQA $\rightarrow$ xbench-DeepSearch, highlighting the convergence of tool synthesis.
Figure 6: EGL dynamics on HLE and accuracy on selected datasets vs. evolution step. The orange curve shows the EGL trend (left axis, scaled by 1000). Blue bars indicate the accuracy (right axis) of agents using HLE toolsets frozen at 10%, 40%, 70%, and 100% of evolution.
5.3 Evolutionary generality loss
To quantify knowledge acquisition, we introduce Evolutionary Generality Loss (EGL), a real-time convergence indicator analogous to training loss. EGL is defined as the cumulative ratio of synthesized tools to invocations:
$\text{EGL}(t)\triangleq\mathcal{L}_{t}=\frac{C_{t}}{U_{t}}=\frac{\sum_{i=1}^{t}c_{i}}{\sum_{i=1}^{t}u_{i}}$ (1)
where $c_{i}$ and $u_{i}$ denote tool synthesis and invocations at step $i$. EGL is inversely correlated with repository generality. As shown in Figure 6 (HLE dataset), initial high values reflect rapid accumulation, which stabilizes asymptotically after $\sim$1,000 queries (40% of the stream).
Figure 7: Evolutionary dynamics across varying batch sizes. (a) Cumulative tool synthesis relative. (b) Mean token consumption per invocation. (c) Aggregate count of tool invocations.
Appendix A provides the complete formal proofs. We first derive the sufficient condition for the monotonic decrease of EGL:
Theorem 5.1 (Descent Condition).
If the marginal synthesis ratio is strictly lower than the cumulative ratio (i.e., $\frac{c_{t+1}}{u_{t+1}}<\frac{C_{t}}{U_{t}}$), the EGL metric decreases ($\mathcal{L}_{t+1}<\mathcal{L}_{t}$).
Thus, EGL declines whenever the marginal creation rate falls below the historical average. Empirically (Figure 6), the trajectory is consistently downward, with transient spikes reflecting domain-specific bursts.
Second, the convergence rate:
Theorem 5.2 (Convergence Rate).
If cumulative tools stabilize ($C_{t}\to C_{\text{max}}$) and invocations grow linearly ($U_{t}\propto t$), EGL converges to zero ($\mathcal{L}_{t}\to 0$) at a rate of $\mathcal{O}(1/t)$.
Figures 5 and 7(c) empirically validate the stabilization of $C_{t}$ and linear growth of $U_{t}$, respectively.
To further validate convergence, we evaluated toolset snapshots (frozen at 10%, 40%, 70%, 100% of the timeline) on 200 stratified queries with further tool synthesis disabled. Figure 6 bar chart reveals a sharp performance inflection between 10%–40%, followed by diminishing returns from 40% to 100%. This saturation confirms the system successfully converges to a high-generality state.
Table 2: Ablation results on 200 queries across varying batch sizes (BS). Metrics include Accuracy (Acc, %), total inference time in minutes (Time), Tool Success Rate (TSR, %), Average Tokens per Tool (ATT), and cost per query (Cost, \$). The condition $\infty$” denotes a Python-interpreter baseline processing queries independently without tool accumulation.
BS Acc Time TSR ATT Cost
[…]
5.4 Ablation study
To investigate the efficacy of tool evolution, we conduct ablation studies using the 200-query EGL subset.
Mechanisms of tool evolution efficacy. We benchmark our approach against a Python-interpreter with the same baseline (GPT-5), representing the asymptotic limit of an infinite batch size ($B\to\infty$) where queries are processed independently without tool accumulation. As detailed in Table 2, this baseline demonstrates markedly inferior performance (40.0) compared to our method (47.5–51.5). Analysis suggests this performance gap stems from the baseline’s lower execution success rate (81.8% vs. >>99%) and significantly higher token consumption (518 tokens per invocation). We posit that the resulting verbosity and frequent execution failures induce contextual contamination, which subsequently degrades reasoning capabilities (see Appendix E).
Table 3: Performance comparison and tool library size of different backend models on DeepSearchQA and FinSearchComp.
Backend DeepSearchQA FinSearchComp
Accuracy # Tools Accuracy # Tools
Gemini 3 Pro 73.5 16 65.0 18
GPT-5 70.6 75 65.2 25
GPT-5-mini 42.5 47 45.5 13
Impact of batch size on parallel evolution. We investigate the influence of batch size $B$ on evolutionary dynamics, evaluating configurations where $B\in\{2,4,8,16\}$ against a Python interpreter baseline ($B\to\infty$). As detailed in Table 2, all batch-evolved agents surpass the baseline, reinforcing the efficacy of tool accumulation. Furthermore, the results indicate that increasing $B$ proportionally reduces processing latency and marginally enhances accuracy without incurring substantial overhead, thereby validating the Parallel Batch Evolution strategy.
The impact of batch size on evolutionary trajectories is further analyzed in Figure 7(a). While larger batches accelerate initial tool crystallization, terminal library sizes remain comparable, suggesting the system converges toward a consistent capability equilibrium independent of $B$. Regarding computational efficiency, Figure 7(b) demonstrates that token expenditures consistently remain below the baseline, decaying toward an asymptotic minimum as the system transitions from costly ab initio tool generation to efficient reuse. A minor anomaly observed for $B=8$ in Table 2—characterized by higher token consumption—is attributed to maximal tool generation and invocation counts (Figure 7(a,c)) resulting from LLM stochasticity. Future research will explore methods to align convergence states across varying batch sizes.
Influence of backend model selection. We evaluate the impact of the underlying foundation model on system performance across the DSQA and FinSearchComp datasets (see Table 3). The results indicate that when instantiated with GPT-5, our agent surpasses the majority of established baselines. Notably, the lightweight GPT-5-mini variant achieves competitive performance, demonstrating significant utility despite its reduced parameter scale. These findings underscore the model-agnostic nature and generalization capabilities of the proposed framework, confirming its robustness across diverse backend architectures and scales.
6 Related work
Self-evolving agents. To transcend the limitations of static designs, general-purpose agent systems must adapt to open-ended, interactive environments in real-time (Gao et al., 2025). Self-evolving agents achieve this by autonomously refining internal components, categorized into workflow optimization, model parameter updates, context management, and tool synthesis. Specifically, MAS-GPT (Ye et al., 2025) and AFlow (Zhang et al., 2025b) generate query-specific workflows, while TT-SI (Acikgoz et al., 2025) and SCA (Zhou et al., 2025) update parameters by generating challenging training problems. Furthermore, ELL (Cai et al., 2025) and Zhang et al. (2025c) evolve by accumulating environmental interaction experience as context. Regarding tool synthesis, LIVE-SWE-agent (Xia et al., 2025) and STELLA (Jin et al., 2025) enable on-the-fly tool creation for software engineering and biomedical research, respectively. Nevertheless, most self-evolving methods necessitate explicit training. Crucially, even approaches closest to our work—those autonomously creating tools at test-time—remain confined to specific domains and lack tool reuse mechanisms, hindering their general applicability.
Tool evolution agents. The paradigm of autonomous tool synthesis has garnered significant attention in recent literature. Pioneering systems such as Voyager (Wang et al., 2023), STELLA (Jin et al., 2025), and LIVE-SWE-agent (Xia et al., 2025) empower agents to generate executable tools for embodied control or software engineering tasks. Notably, analyses of the tool embedding space in LIVE-SWE-agent (Xia et al., 2025) reveal distinct clustering patterns among functional equivalents; this observation directly informed the design of our absorbing mechanism, which utilizes clustering to consolidate redundant utilities. Regarding tool reuse, Alita-G (Qiu et al., 2025) facilitates the persistence of successful tools, though its optimization trajectory is predicated on ground-truth supervision. In the context of inference-time adaptation, Lu et al. (2026) proposed a test-time tool evolution framework governed by a cost-sensitive objective function, yet its application remains narrowly scoped to scientific reasoning. Collectively, these approaches prioritize the optimization of existing assets or domain-specific synthesis, distinguishing them from our framework’s focus on self-evolution in open-ended environments.
Agent systems for open-ended tasks. Recent research has sought to construct general-purpose agents capable of cross-domain operation. DeepAgent (Li et al., 2025) utilizes a scalable toolset to address diverse tasks but lacks the capacity for autonomous environmental adaptation, leading to performance limitations. Agent KB (Tang et al., 2025) attempts to solve cross-domain problems by summarizing execution experiences from other agents; however, its performance is constrained by a deficiency in necessary tools. In contrast, our approach achieves generality by continuously creating, reusing, and optimizing tools in-situ, effectively overcoming the limitations of static toolsets and experience-based adaptation.
7 Discussion and future work
While this study establishes the efficacy of in-situ self-evolution through tool synthesis, several avenues for future research merit rigorous inquiry to fully realize the potential of autonomous agents.
Paradigm parallels: towards system-level pre-training for agentic systems. The seamless transition and performance consistency observed between our zero-start and warm-start settings offer more than just empirical validation; they suggest a fundamental paradigm shift. The clear convergence curves of the tool library indicate that “task-solving capability” is not merely a collection of ad-hoc heuristics, but a generalizable pattern that can be learned and distilled.
This implies that the field is approaching a “pre-training and post-training” era for agentic systems, mirroring the trajectory of LLMs. We envision a future where multi-agent systems undergo system-level pre-training on massive, broad-spectrum task datasets. This process would allow agents to distill a “foundation toolset”—a set of converged, generalizable primitives—before deployment. Consequently, such pre-trained agents would possess intrinsic generalization capabilities, enabling them to tackle novel downstream tasks largely through the composition of existing reliable tools, thereby minimizing or even eliminating the need for expensive test-time evolution. Formalizing the methodologies for this “agentic pre-training” is a critical priority for future research.
Co-evolution of memory and workflow. Currently, our framework validates the self-evolving paradigm primarily through tool generation. However, tools constitute a necessary but insufficient condition for scenarios demanding high personalization or complex process management. For instance, personalized assistants require persistent, evolving memory structures to align with user preferences, while intricate tasks (e.g., deep research) necessitate structured workflow evolution—such as generating bespoke planning protocols for multi-agent coordination. A critical direction is extending the evolutionary mechanism to encompass the co-evolution of memory architectures and workflow policies, enabling the system to adapt its internal state and execution logic alongside its functional capabilities.
Evolutionary stability and regularization. The stability of tool evolution warrants further investigation. Due to the inherent stochasticity of LLM generation, toolsets can exhibit variance across experimental runs, as evidenced by our batch size ablation. Ensuring the consistent convergence of the tool library is vital for system reliability. Future work will focus on developing regularization strategies to guarantee the determinism of the evolutionary process in open-ended environments.
Optimization of parallel batch evolution. Finally, the nuances of the batch evolution strategy present rich opportunities for optimization:
  • Curriculum learning effects: As shown in Figure 5, the sequence of incoming queries significantly influences the convergence trajectory; delaying queries that require foundational primitives can impede system maturation. Investigating optimal query ordering is a key direction.
  • Intra-batch diversity trade-offs: Intra-batch diversity presents a complex dynamic. Low diversity allows the absorbing mechanism to function as a form of Best-of-N test-time scaling, leveraging redundancy to select the optimal implementation—a contrast to traditional gradient training where high batch diversity is preferred. Balancing this quality-assurance benefit against the efficiency of capability evolution is crucial.
  • Adaptive scheduling: Dynamic batch sizing offers a promising avenue for optimization. Smaller batch sizes in early stages could foster the robust consolidation of general knowledge via redundancy, while larger batches in the post-convergence phase could maximize throughput for corner cases. Developing autonomous agents capable of dynamic batch scheduling based on convergence signals remains a worthy research direction.
8 Conclusion
In this work, we presented the In-Situ Self-Evolving framework, enabling LLM-based agents to autonomously adapt and evolve within open-ended environments. Through the Yunjue Agent and the proposed Parallel Batch Evolution strategy, we demonstrated that treating tools as dynamic vehicles for knowledge crystallization allows for robust zero-start learning and the emergence of generalized, transferable capabilities. Our empirical results validate that this approach achieves state-of-the-art performance across heterogeneous benchmarks. Furthermore, warm-start experiments confirm that the accumulated generalized knowledge can be seamlessly transferred to novel domains, enabling the agent to adapt to shifting environments. We release our codebase, evaluation and system traces to the open-source community to facilitate further research. Broadly, future work will aim to unify tool evolution with memory and workflow adaptation, advancing toward more autonomous and general-purpose agentic systems.
Appendix A Mathematical proof
This section details the formal proofs for Theorems 5.1 and 5.2, originally presented in Section 5.3.
A.1 Proof of theorem 5.1
We begin by establishing the sufficient condition for the monotonic decrease of the EGL metric at step $t+1$ (i.e., $\mathcal{L}_{t+1}<\mathcal{L}_{t}$). By definition:
$\mathcal{L}_{t+1}=\frac{C_{t}+c_{t+1}}{U_{t}+u_{t+1}}<\frac{C_{t}}{U_{t}}=\mathcal{L}_{t}$ (2)
Since the total invocation count $U_{t}$ is strictly positive, we cross-multiply without reversing the inequality:
$(C_{t}+c_{t+1})\cdot U_{t}<C_{t}\cdot(U_{t}+u_{t+1})$ (3)
Expanding the terms yields:
$C_{t}U_{t}+c_{t+1}U_{t}<C_{t}U_{t}+C_{t}u_{t+1}$ (4)
Subtracting the common term $C_{t}U_{t}$ and rearranging results in the Descent Condition:
$\frac{c_{t+1}}{u_{t+1}}<\frac{C_{t}}{U_{t}}$ (5)
This derivation demonstrates that $\mathcal{L}_{t}$ decreases if and only if the marginal ratio of synthesis to invocation at the current step ($c_{t+1}/u_{t+1}$) is strictly lower than the historical cumulative ratio ($C_{t}/U_{t}$). During early evolutionary stages, extensive exploration ($c_{t+1}>0$) may induce fluctuations. However, as the agent transitions to an exploitation phase ($c_{t+1}\to 0$ while $u_{t+1}\geq 1$), the marginal ratio vanishes, guaranteeing satisfaction of the inequality and ensuring the strict decrease of the metric.
A.2 Proof of theorem 5.2
To establish the asymptotic dynamics of $\mathcal{L}_{t}$, we introduce the following assumptions governing the evolutionary process:
Assumption A.1 (Knowledge Boundedness).
The task domain admits a finite set of optimal primitives sufficient to resolve all potential queries. As the agent evolves, the cumulative count of synthesized tools $C_{t}$ converges to a finite upper bound $C_{\text{max}}$. Consequently, marginal tool generation vanishes asymptotically:
$\lim_{t\to\infty}C_{t}=C_{\text{max}}\implies\lim_{t\to\infty}c_{t}=0$ (6)
We posit this as a mild condition. Empirically, as validated in Figure 6, the marginal growth of synthesized tools diminishes toward zero once the agent consolidates sufficient generalized knowledge.
Assumption A.2 (Continuous Activity).
We assume every newly synthesized tool is immediately invoked at least once, ensuring $0\leq c_{t}\leq u_{t}$ and $0\leq\frac{c_{t}}{u_{t}}\leq 1$. Operationally, if a synthesized tool remains uninvoked, it is discarded post-processing to satisfy this constraint. Additionally, we assume every query necessitates at least one tool invocation ($u_{t}\geq 1$), with the invocation count fluctuating around a stationary mean $\mu\geq 1$. This implies that cumulative invocations $U_{t}$ grow linearly with respect to $t$:
$U_{t}\approx\mu\cdot t\quad\text{as }t\to\infty$ (7)
Note that for the boundary condition $u_{t}=0$, we define $c_{t}/u_{t}=0$ to avoid singularity. Figure 7(c) provides empirical evidence supporting the linear growth of $U_{t}$ throughout the evolutionary steps.
Based on Assumptions A.1 and A.2, we analyze the limit of $\mathcal{L}_{t}$ as $t\to\infty$:
$\lim_{t\to\infty}\mathcal{L}_{t}=\frac{\lim_{t\to\infty}C_{t}}{\lim_{t\to\infty}U_{t}}\approx\lim_{t\to\infty}\frac{C_{\text{max}}}{\mu\cdot t}$ (8)
Since $C_{\text{max}}$ and $\mu$ are positive constants:
$\lim_{t\to\infty}\mathcal{L}_{t}=0$ (9)
This confirms that EGL functions mathematically as a loss metric, converging to zero at a rate of $\mathcal{O}(1/t)$ as system stability is achieved.
Appendix B Algorithm
Algorithm 1 Yunjue Agent Execution and Evolution Pipeline
1:Input: Query Batch $\mathcal{Q}_{t}=\{q_{t,1},\dots,q_{t,B}\}$, Global Tool Library $\mathcal{T}_{t-1}$
2:Output: Answers $\mathcal{A}_{t}$, Updated Library $\mathcal{T}_{t}$
[…]
4:for each query $q_{t,i}\in\mathcal{Q}_{t}$ in parallel do
5:  $\mathcal{R}_{miss}\leftarrow\emptyset$, $\text{done}\leftarrow\textbf{false}$
6:  while not done do
7:   $\omega_{req}\leftarrow\textbf{Manager}(q_{t,i},\mathcal{T}_{t-1},\mathcal{R}_{miss})$ {Analysis required tools}
8:   if new tool required then
9:    $\mathcal{P}_{t,i}\leftarrow\textbf{ToolDeveloper}(q_{t,i},\omega_{req})$ {Synthesize new tools}
10:    $\mathcal{T}_{local}\leftarrow\mathcal{T}_{t-1}\cup\mathcal{P}_{t,i}$
11:    $\mathcal{T}_{new}\leftarrow\mathcal{T}_{new}\cup\mathcal{P}_{t,i}$
12:   else
13:    $\mathcal{T}_{local}\leftarrow\mathcal{T}_{t-1}$
14:   end if
15:   $r_{t,i},\mathcal{R}_{miss}\leftarrow\textbf{Executor}(q_{t,i},\mathcal{T}_{local})$ {Execute and update capability missing report}
16:   if $\mathcal{R}_{miss}=\emptyset$ then
17:    $\text{done}\leftarrow\textbf{true}$
18:   end if
19:  end while
20:  $a_{t,i}\leftarrow\textbf{Integrator}(q_{t,i},r_{t,i})$ {Integrate final answer}
21:  $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t}\cup\{a_{t,i}\}$
22:end for
[…]
25:return $\mathcal{A}_{t}$, $\mathcal{T}_{t}$
Appendix C Experiment setup
Table 4: Statistics of the datasets.
Dataset Domain Language # Queries
HLE General reasoning EN 2,500
DeepSearchQA General QA EN 900
FinSearchComp (T1 & T2) Financial search EN & CH 391
xBench-ScienceQA General science CH 100
xBench-DeepSearch General research CH 100
Figure 8: EGL dynamics on the DSQA and FSC datasets.
Data statistics are summarized in Table 4.
Evaluation protocols. We follow the standard evaluation protocols released with each dataset:
Baselines sources. We collect baseline results from their official reports or leaderboards. For HLE, results for Gemini Deep Research and Gemini 3 Pro are from (Google, 2025); MiroThinker-v1.0-72B and Tongyi DeepResearch are from (Team et al., 2025); GPT-5.2 Pro is from (OpenAI, 2025); Seed-1.8 is from (ByteDance, 2025); Claude 4.5 Opus is from (Anthropic, 2025); and Co-Sight is from (Zhang et al., 2025a). DeepSearchQA results are from (Gupta et al., 2025). FinSearchComp results are from (ByteDance, 2025). For xBench, results on ScienceQA and DeepSearch are retrieved from (xBench Team, 2025b) and (xBench Team, 2025a), respectively. Additionally, for xBench-DeepSearch, we include M-GRPO (Hong et al., 2025) and AWorld (Yu et al., 2025).
Implementation details. To ensure experimental consistency, all nodes in our system are instantiated using the same backend LLM: Gemini 3 Pro. The batch size is set to $B=16$ for all datasets. The Tool Developer module is explicitly powered by Codex (Zaremba et al., 2021) to facilitate robust code generation. We main a fixed temperature of 0.7 across all LLM invocations. To simplify the use of multimodal capabilities, we encapsulate image processing functionality into a dedicated tool, ensuring utilization of the same backend LLM. Our agent system is built based on LangGraph (LangChain, 2024). All prompts in our system adopt the Markdown template format supported by Jinja (The Pallets Projects, 2007).
Appendix D More evaluation results
Table 5: Experimental results across five benchmarks of the Yunjue Agent in the zero-start setting. “Token” denotes the average token consumption per query, “Exec. Ratio” indicates the proportion of tokens consumed by the Executor, and “Cost” denotes the monetary expense based on token usage.
Dataset Accuracy (%) Token Exec. Ratio (%) Cost (\$)
HLE $48.0_{\pm 0.2}$ 421k 95.1 1.48
DeepSearchQA $73.5_{\pm 0.4}$ 718k 95.6 1.78
FinSearchComp $65.0_{\pm 0.1}$ 474k 94.5 1.30
xbench-ScienceQA $76.5_{\pm 0.6}$ 207k 94.2 0.73
xbench-DeepSearch $59.7_{\pm 0.2}$ 569k 94.7 1.48
Detailed results on five benchmarks. Table 5 summarizes the detailed performance of the Yunjue Agent in the zero-start setting across five datasets, where accuracy is calculated as the mean and standard deviation over three runs using the LLM-as-judge.
More results on evolutionary generality loss (EGL). Figure 8 tracks the EGL trajectories across the DSQA and FSC benchmarks, while Figure 9 dissects the impact of batch size on optimization behavior. Collectively, these visualizations validate the asymptotic stability of our approach across varying experimental conditions.
Figure 9: EGL dynamics under different batch sizes.
Technical insights. We observed that Gemini 3 Pro possesses strong reasoning capabilities and internal knowledge, often exhibiting a sense of “confidence”, which is manifested by its tendency to rely on fewer tools to complete tasks. However, it frequently suffers from hallucinations and often exhibits weaker instruction-following capabilities, necessitating iterative prompt engineering. In contrast, GPT-5 and GPT-5-mini often appear more “cautious”, typically requiring the invocation of more tools, planning, and reasoning steps to iteratively verify task results. This leads to increased tool creation and longer execution histories. This phenomenon is particularly pronounced in GPT-5 (as shown in Table 3, the GPT-5 series clearly creates more tools). Nevertheless, the GPT-5 series demonstrates exceptional instruction-following abilities; adding constraints to their prompts generally yields behavior consistent with expectations.
Figure 10: Yunjue Agent execution pipeline. Upon receiving a query, the agent autonomously synthesizes, deploys, and iteratively refines tools to derive the final response, seamlessly integrating generation and execution.
Figure 11: Illustration of tool absorbing mechanism. Following batch execution, functionally analogous tools are identified via clustering and consolidated into a generalized, compact toolset, effectively pruning redundancy.
Figure 12: Python-only baseline failure mode. Unlike our approach, the baseline indiscriminately accumulates raw execution traces—including erroneous attempts—within the context window, leading to severe contextual contamination and reasoning degradation.
Appendix E Case study
Tool evolution. Figure 10 illustrates the end-to-end execution pipeline of the Yunjue Agent. When presented with a query, the Manager delegates tool creation to the Tool Developer. The agent effectively instantiates necessary tools and, crucially, demonstrates self-correction capabilities. When an initial tool fails (e.g., a PDF reader failure), the system detects the error, refines its strategy, and successfully derives the final answer, highlighting the seamless integration of tool generation, execution, and refinement.
Knowledge consolidation via absorbing. Figure 11 depicts the tool absorbing mechanism. Following batch execution, the initial tool pool often contains redundant or highly specific tools. The system employs clustering to identify functionally analogous tools and merges them into a compact, generalized toolset (e.g., consolidating multiple web scrapers into a single robust tool). This process effectively prunes redundancy, ensuring the long-term efficiency and generalizability of the tool library.
Comparison with Python-only baseline. Figure 12 contrasts our approach with a typical Python-only baseline failure mode. In complex tasks like medical diagnosis, the baseline often struggles with library dependencies and engages in high-friction debugging loops. Unlike our approach, the baseline accumulates raw, erroneous execution traces within its context window, leading to severe context contamination and reasoning degradation.
Tool similarity and clustering To verify the semantic alignment of tools between independent evolutionary runs (as visualized in Figure 4), we present a direct code comparison. We contrast send_http_request, under a zero-start setting, with fetch_api_data, under a warm-start setting, evolved for DeepSearchQA. Despite distinct origins, both converge on core HTTP functionality, necessitating similar dependencies and logic structures.
Tool: send_http_request Tool: fetch_api_data
Appendix F Prompts
Manager Tool Developer Executor Integrator Aggregator Merger